OpenStudy (danmac0710):

A tricky little modulus (absolute value) equation. Any ideas please?: abs(2x+3) = 3 * abs(x) -1

OpenStudy (amistre64):

looks simple enough ...

OpenStudy (amistre64):

where are your vertexes located?

OpenStudy (amistre64):

if you can determone your vertexes, you can redefine the abs function into its parts and just focus on the intersection of 2 lines 

OpenStudy (danmac0710):

Thanks for that hint.... working on it!

OpenStudy (danmac0710):

Do you always do these ones graphically?

OpenStudy (amistre64):

its prolly not necessary, but i work better with a visual ... you could just divvy it up into intervals to test the 4 line equations that make it up ...

OpenStudy (amistre64):

from -inf to -3/2; we play with: -(2x+3) = -3x-1 from -3/2 to 0; we play with: (2x+3) = -3x-1 from 0 to inf; we play with: (2x+3) = 3x-1

OpenStudy (amistre64):

spose its only 3 setups not 4

OpenStudy (danmac0710):

cheers man. That's helpful. Never done these type before.

OpenStudy (phi):

that is a nice visual. if you do algebra, you have 4 cases: drop the abs value on both expressions + + negate the left expression, - + negate the right expression + - and negate both - - from the picture we see we will get 2 "non-solutions"